OpenStudy (anonymous):

What is the sum of the first five terms of the geometric sequence in which a1 = 10 and r = 1/2? Express your answer as an improper fraction.

OpenStudy (johnweldon1993):

The sum of the first 'n' terms can be given as.. \[\large a_n = a\frac{1 - r^{n}}{1 - r}\] so for the first 5 terms here...with a = 10 and r = 1/2 we have... \[\large a_5 = 10\frac{1 - \frac{1}{2}^{5}}{1 - \frac{1}{2}}\] Can you solve this?

OpenStudy (anonymous):

well for 1-1/2^5 I got 9.6, would I multiply that part by 10? I'm confused @johnweldon1993

OpenStudy (johnweldon1993):

Well 9.6875...hard to round that but I would just use this... so yes multiply that by 10 to get 9.6875 so now you have... \[\large S_5 = \frac{9.6875}{.5}\] Right? Now just evaluate that...*or actually your question states to leave your answer in an improper fraction...so this may be all you need to do...

OpenStudy (anonymous):

so I dont need to simplify that? the sum of the 5 numbers would simply be 9.6875/.5?

OpenStudy (johnweldon1993):

If you want to simplify it go ahead...but based on what the question states...this would be the correct form... if you want to evaluate it obviously it would be 19.375

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

it says that the correct answer was 155/8

OpenStudy (johnweldon1993):

Yeah that is the correct answer....both in numerical form = 19.375 We just had the wrong fraction apparently..still correct though